OpenStudy (anonymous):

Solve the equation for a 3a^2+22a+35=0

OpenStudy (apoorvk):

factorise!!!! split the middle term!!!! like this---> 3a^2 + 15a + 7a + 35 = 0 Now can you find common terms and take it further?

OpenStudy (anonymous):

Your equation is in the form of ax^2 + bx + c = 0 Use the formula \[-b \pm \sqrt{b ^{2}-4ac}/ 2a\]

OpenStudy (jlastino):

Ok looking at the equation you can solve it by factoring, like Apoorvk said :D

OpenStudy (anonymous):

11/210 or 5/210